OpenStudy (anonymous):

A particle of mass m moves with momentum of magnitude p. (a) Show that the kinetic energy of the particle is K = p2/(2m) (b) Express the magnitude of the particle's momentum in terms of its kinetic energy and mass.

OpenStudy (anonymous):

P=mv squaring.. P^2=m^2.v^2 dividing by m P^2/m=mv^2 multiplying by 1/2 P^2/2m=K.E.